OpenStudy (anonymous):

doing process of elimination need the solution to; 4A+4B=24 -4A+B=21 PLEASE HELP

OpenStudy (anonymous):

add the two equations together and solve for B

OpenStudy (anonymous):

DID THAT ENDED UP WITH 9

OpenStudy (anonymous):

5B=45 so B=9, now use either equation to get A.

OpenStudy (anonymous):

substitute B=9 in the equation then solve for A

OpenStudy (anonymous):

THIS IS WHERE I GET LOST I CANT FIGURE IT OUT FROM THERE????

OpenStudy (anonymous):

-4A+B=21 then sub B=9 , -4A+9=21...

OpenStudy (anonymous):

A IS -3 THEN????

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

-4A=21-9, -4A=12,A=12/-4,A=-3. Yippy

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

SO IN MY SOLUTION WOULD BE (-3,15/4) IS THAT RIGHT??

OpenStudy (anonymous):

dunno

OpenStudy (anonymous):

Too check your answer, put A=-3,B=9 into both equations, -4A+B=21, -4(-3)+9=21,21=21. The last equation is an identity statement. It is a "true" statement. So you just proved that A=-3,B=9 is a solution of this equation. Next you need to do the same for the other equation. It MUST also work there.

OpenStudy (anonymous):

is this schoolwork????????????????????

OpenStudy (anonymous):

The solution of a system of equations is the set (group) of all pairs (A,B) that satisfy the system. (what I did above is how you show that the system is satisfied). so your answer is the pair (A value, B value)

OpenStudy (anonymous):

OH thank you so much now that you have explained the second part i get it!!!!